Tohban Report 2013-01-09

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 2 Jan 2013
End Date: 9 Jan 2013
Tohban: Juan Carlos Martinez Oliveros
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

Solar activity was moderate this week. There are currently 14 active regions on disk with active regions. AR 11654 is showing some activity, and is at the moment on the East limb. Activity is expected to remain low-moderate with a chance for an M class flare in the coming days.

How many GOES flares occurred?

 Flares above B, C, M, X class were      8    34     1     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      7    17     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      4     8     0     0

There were RHESSI flares/GOES flares 213 / 43 over the time range 02-Jan-13 09-Jan-13

Memory Management

The SSR was emptying at the en of the daily passes during the beginning of the week. After operations the SSR fill level was between 7 and 25%.

Spacecraft Status

The cold plate 2 is at 107 respectively. Cryopower at about 70.6 Watts.

Spin rate is at 14.8

Data Gaps

There are several VC1 and VC3 data gaps reported, however all the data is on the ground. It is likely that the data was not processed due to a malfunction of the Apollo server and are currently being processed.

Due to a problem with ITOS on January 5 there is a VC1 data gap. The ITOS instance that supports BGS passes hung up. As a result, no commands were sent from ITOS to the spacecraft to initiate VC1 dumps with any of the BGS passes. No gaps for the VC3 data were reported, as these data is transmitted by onboard (ATS) commands.

N_GAPS           1 
GAP START TIME              GAP END TIME                   GAP (SEC) 
2013-01-05T16:20:00.000 -- 2013-01-06T00:00:00.000       27600.000 
N_GAPS           1 
GAP START TIME              GAP END TIME                   GAP (SEC) 
2013-01-05T16:20:00.000 -- 2013-01-06T00:00:00.000       27600.000 

Detector issues

On January 3 events on the rear detectors were during spacecraft eclipse. These events started being recorded at 2013-004:00:00 UTC.

On January 8 at 07:26:26 UTC the high voltage of detector 2 was raised by five steps (206 -> 211). There was a modest increase in front slow and reset rates, but no sign of segmentation. In the following days a slight increase of the noise was observed. The voltage is now at 4117 V.

Other notes

The decimation state was changed from active vigorous to normal vigorous on January 3 at 16:30:38 UTC

Spacecraft Management

Decimation Normal/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) Taking data at nights
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? No
Personal tools